OpenStudy (anonymous):

Write the equation of a quadratic function with vertex at (-1,1) f(x) = (x + 1)^2 + 1 f(x) = (x + 1)^2 - 1 f(x) = (x - 1)^2 + 1 f(x) = (x - 1)^2 - 1

OpenStudy (anonymous):

If the vertex is (h, k), then the equation is: y = a(x - h)^2 + k Here, you can set a=1 and your h= -1 and k = 1. So, just substitute at this point.

OpenStudy (anonymous):

huh..

OpenStudy (anonymous):

Since you are letting a=1, the equation is: y = (x - h)^2 + k So, now, you can ignore "a". So now just put your values for h and k in. What do you get so far? Show your work.

OpenStudy (anonymous):

By substituting, that means where you see h and k in the equation, put the value of h and k in the equation instead.

OpenStudy (anonymous):

so its f(x) = (x - 1)^2 + 1 ?

OpenStudy (anonymous):

For the part within the parentheses, you have [x - (-1)] or (x - -1) which becomes (x + 1)

OpenStudy (anonymous):

so its this then f(x) = (x + 1)^2+ 1

OpenStudy (anonymous):

Yes! Good work, you got it!

OpenStudy (anonymous):

Good luck in all of your studies and thx for the recognition! @kimdre10